2. For the following Cobb-Douglas production function, q=f(L,K) = _0.45 0.7 a. Derive expressions for marginal product of labSHOW ALL WORK!!!
Answer #1

q f(L,K) = 10.45 Kon a.) MP = df (LK) 0.45 Loss Kon MPK = df (L,K) 0.7 10.45 K-0.3 6.) MRTS = MP MP MPK = 0.45 (-0.55 KO7 0.7d.) / (1.51, 1.5k) (1.5) is f(L,K) 11.59 & CLK) Output increases by approximately 5.9% if the firm increased each input by s